# The Philosophical Foundations of Growing AI Like A Child

### v1

@@ 4 Interpreting Core Knowledge Deficits in LLMs
Why do LLMs fail to acquire core knowledge through scaling, despite having access to vast linguistic datasets spanning virtually the entire internet and the computational power to process them? The answer lies in the fundamental differences between LLMs and humans—the only known intelligent agents that possess core knowledge. By systematically examining these differences in architecture and learning processes, we can develop testable hypotheses to identify the specific reasons behind LLMs’ core knowledge deficits. If future models are designed based on these hypotheses and their performance begins to align with human abilities, it would provide a proof of concept for addressing these limitations. The following sections explore three key factors that may contribute to core knowledge deficits in LLMs. Below, we explore three potential factors that may underlie core knowledge deficits in LLMs.
@@ 4.1 Lack of Hardwired Domain-specific Faculties
To begin with, in line with a long tradition in nativist epistemology (Locke, 1824; Chomsky, 1980; Cowie, 2002), it has been suggested that fundamental knowledge of the world—such as core knowledge—must be built upon ”hardwired” cognitive faculties that are innate to the brain at birth. Neurobiological findings indicating specialized modularity in the human brain for processing certain types of information, such as numbers, objects, and social relations, support this perspective (Spelke & Kinzler, 2007). If this view holds, LLMs’ core knowledge deficits may be fundamentally unresolvable, as their foundational architecture lacks such hardwired domain-specific vehicles.
@@
However, this argument is inconclusive at best. To start, empirically validating the existence of innate, domain-specific cognitive mechanisms in humans is challenging, as what appears to be ”innate knowledge” could instead be ”early knowledge”—the result of domain-general learning mechanisms rapidly self-organizing into primitive forms of functional modularity upon first exposure to the environment during infancy, as Piaget posited in terms of accommodation (Piaget, 1952). Moreover, even if humans do possess built-in cognitive faculties, they may not be strictly necessary for any system capable of acquiring core knowledge. To claim otherwise would impose a learnability constraint, asserting that certain cognitive capacities cannot be learned from data using domain-general empiricist mechanisms and instead require innate structures (Long, 2024).
Ultimately, a key argument regarding the learnability of core knowledge is that infants lack explicit symbolic systems and that there are no widely accepted mechanisms by which pre-linguistic intelligent systems could form representations of domain-specific concepts solely through perceptual interaction (Carey, 2011). However, given LLMs’ strong ability to manipulate linguistic symbols, they may be able to leverage abstract concepts through domain-general learning mechanisms to acquire core knowledge representations (Long, 2024).
@@ 4.2 Buried Too Deeply
Secondly, it may be argued that LLMs may already possess core knowledge but are unable to extract and apply it in reasoning. Since the early days of connectionist research, it has been suggested that artificial neural networks, due to their reliance on distributed representations, inevitably face a challenge: as network size increases through scaling, retrieving knowledge representations for reasoning becomes increasingly difficult and computationally costly (Hinton et al., 1986; Clark, 1992). As Chalmers (1990) remarked:
@@
This concern is particularly relevant for core knowledge. High-level details required for complex tasks are relatively easy to retrieve, as their representational patterns are likely clustered within the network, such as those encoding a specific historical event or a recipe for a particular dish. In contrast, identifying and utilizing core concepts in simple tasks is significantly more challenging for LLMs, as these concepts are distributed across numerous samples in the dataset with roughly the same level of salience. For example, while a vast number of instances in a model’s training data may implicitly demonstrate perceptual constancy—the principle that an object’s identity remains unchanged despite variations in sensory representation—this concept is not explicitly reinforced in a way that facilitates reliable retrieval (Garrigan & Kellman, 2008; Green, 2024). Since basic concepts like this is embedded in diverse contexts rather than tied to specific, distinguishable patterns, LLMs may struggle to isolate and apply it systematically, leading to inconsistent or surface-level reasoning when tested on related tasks. Consequently, while scaling may provide sufficient computational power and data to encode core knowledge, models remain unable to effectively employ it for reasoning. This is because its compositional structure is encoded in an excessively dispersed manner across the vast parameter space of the network (Shani et al., 2023).
4.3 Groundless Learning Process
Finally, a key distinction between human and machine learning lies in the temporal dynamics of data exposure. Humans progress through a structured developmental trajectory, where cognitive and representational capacities are initially limited, allowing them to process information only within a constrained scope. As they mature, their cognitive system gradually expands, building upon foundational core knowledge to integrate increasingly complex abstractions. This incremental learning process enables humans to develop a deep, structured understanding of the world, where high-level reasoning emerges as a natural extension of foundational cognitive abilities (Pezzulo et al., 2013).
In contrast, LLMs do not follow this grounded developmental process. Instead, they are trained from the outset with access to vast datasets encompassing a broad spectrum of human knowledge, including highly abstract and complex information. Unlike humans, who first acquire intuitive principles through direct sensorimotor experiences before developing abstract knowledge on top, LLMs process high-level concepts alongside low-level ones without any structured progression. This lack of developmental scaffolding means that while LLMs can generate responses that resemble advanced reasoning, they may lack the fundamental conceptual grounding that allows humans to apply knowledge flexibly and coherently across different contexts (Mitchell & Krakauer, 2023).
However, this difference in developmental trajectory does not necessarily preclude LLMs from acquiring core knowledge. If trained on data similar to those available to a child, LLMs might develop a structured understanding of fundamental concepts in a way that mirrors human learning. Rather than relying solely on large-scale text-based training, models could benefit from multimodal learning, where training data prioritizes direct representations of low-level cognitive concepts through rich perceptual information. By leveraging their ability to process symbolic representations, LLMs may be able to construct conceptual frameworks akin to core knowledge developed by humans.
@@ 5 Moving Forward: Growing AI Like A Child
The interpretations of potential reasons for core knowledge deficits in LLMs suggest a testable hypothesis: LLMs may be capable of acquiring core knowledge if trained on environmental stimuli similar to those available to a child. By leveraging their ability to process symbolic representations, LLMs might develop a structured understanding of fundamental concepts upon similar external stimuli accessible to a child developing the same concepts.
For this proposal to be viable, several key conditions would have to be considered:
@@
    First, sensory grounding-models must be trained using multimodal inputs, incorporating both visual and linguistic data to approximate the rich perceptual experience of early human learning.
@@
    Second, large-scale exposure—a sufficiently vast dataset is needed to ensure that core knowledge representations are robust and not lost due to extraction difficulties inherent in distributed neural representations.
@@
    Third, minimizing environmental confounds—unlike a child, LLMs have unrestricted access to abstract information in their training data, making them prone to shortcut-taking by exploiting spurious correlations rather than developing genuine conceptual understanding. Ensuring that training data is curated to prevent such biases will be crucial.
To this end, this paper outlines a practical solution that meets these conditions: the large-scale development of synthetic data using physical engine, operate on a novel cognitive prototyping strategy.
@@ 5.1 Cognitive Prototype
To generate training datasets that enhance the saliency of direct representations of low-level cognitive concepts while minimizing environmental confounds, a promising approach is to directly train MLLMs on cognitive experiments commonly used in developmental psychology laboratories. The recent adaptation of these experiments into machine-readable formats, as seen in cognitively inspired benchmarks (Binz & Schulz, 2023; Li et al., 2025), provides a viable framework for constructing such datasets.
@@
Each low-level cognitive concept can manifest in multiple ways, necessitating diverse experimental paradigms for its assessment. To systematically structure these training datasets, we propose the cognitive prototype, a standardized framework for curating data for each cognitive ability. A cognitive prototype consists of detailed specifications of a cognitive experiment that operationalize the target concept, along with schematic descriptions of task conditions that allow for controlled variations. This structured approach enables systematic data generation at scale while preserving the conceptual rigor of experimental paradigms.
For example, the Three Mountain Task can be used to construct a cognitive prototype for training perspective-taking. In its standard form, this task involves presenting a child with a model featuring three distinct mountains—one covered in snow, another marked by a red cross, and the third topped with a hut. After allowing the child to observe the model from all angles, an experimenter introduces another individual, represented by a doll who views the model from a different vantage point. The child is then shown a set of photographs depicting various perspectives of the model and is asked to identify which image accurately represents what the other person sees (Piaget & Inhelder, 1969). Once the experimental paradigms are specified within the cognitive prototype, data generation can be systematically conducted at a large scale by varying task-relevant conditions in the experiments. These variations may include adjustments in the placement and appearance of the doll, changes in the observer’s viewpoint, modifications to the model’s design, or alterations in the probing questions (Gao et al., 2024). This approach ensures diverse and representative training data that reinforces the foundational principles of core knowledge while preventing models from relying on spurious correlations. By leveraging cognitive prototypes for structured data generation, this strategy offers a scalable and theoretically grounded approach to training MLLMs on core cognitive abilities, addressing key limitations in their current developmental trajectory.
@@ 5.2 Synthetic Data Via Physical Simulations
Once the cognitive prototypes are established, systematically varying task conditions enables the generation of a vast dataset, where each instance presents a unique composition of task elements while consistently exemplifying the same underlying cognitive construct. However, manually labeling such large-scale data would be prohibitively costly and significantly limit the dataset’s scope. To address this challenge, we propose the use of synthetic data generation, leveraging computational tools to create large, diverse, and precisely labeled datasets.
Rather than relying on real-world examples, data can be generated using physics engines such as Mujoco (Todorov et al., 2012) and the upcoming Genesis (Zhou et al., ). These engines allow for precise control over task parameters, enabling the automated creation of experimental scenarios that adhere to cognitive science paradigms while eliminating the inconsistencies and biases present in real-world datasets. By explicitly defining the variables within each experimental paradigm and establishing structured protocols for administering variations, physics engines can systematically produce a vast array of training instances. This ensures that each generated example accurately operationalizes the target cognitive concept while introducing controlled diversity in conditions. Taken together, this approach offers a scalable and theoretically grounded method for training MLLMs on core knowledge. By combining cognitive prototypes with physics-based synthetic data generation, we can create high-quality datasets that reinforce fundamental cognitive abilities, reduce reliance on spurious correlations, and ultimately provide a more structured developmental foundation for machine intelligence.

### v2

@@ 4 Interpreting Core Knowledge Deficits in LLMs
Why do large language models (LLMs) fail to acquire core knowledge through scale alone, despite access to vast linguistic datasets and the computational resources to process them? A promising approach to this question lies in systematically comparing foundation models with humans, the only known agents that reliably exhibit core knowledge. By analyzing differences in architectural inductive biases, learning dynamics, and modalities of experience, we can formulate testable hypotheses to explain LLMs’ limitations in this domain. These hypotheses can, in turn, inform the design of future models. If such models demonstrate improved alignment with human-like core knowledge abilities, it would serve as a proof of concept that these limitations are not intrinsic to LLMs, but instead stem from specific design choices or training regimes. Below, we outline three potential interpretations of the reasons underlying core knowledge deficits in MLLMs.
@@ 4.1 Lack of Hardwired Domain-specific Faculties
To begin with, drawing from a longstanding tradition in nativist epistemology (Kant, 1781; Locke, 1824; Chomsky, 1980; Cowie, 2002), it has been argued that core knowledge must be grounded in innate, domain-specific cognitive faculties—neural structures that are biologically “hardwired” at birth. Neurobiological evidence indicating the eixstence of specialized cortical regions for processing particular domains such as numerosity, object individuation, and social agency have been cited as supports for this view (Spelke & Kinzler, 2007). From this perspective, core knowledge deficits in LLMs may be fundamentally unresolvable, given that their architectures lack such dedicated, domain-specific inductive biases.
@@
However, this argument remains inconclusive at best. First, empirically demonstrating the existence of innate domain-specific mechanisms in humans is notoriously difficult. What appears to be ”innate” may instead reflect ”early knowledge”, namely structured patterns that arise from domain-general learning mechanisms rapidly self-organizing into functionally modular systems upon early environmental exposure, consistent with Piaget’s notion of accommodation (Piaget, 1952). Moreover, even if certain faculties are biologically innate in humans, they may not be computationally necessary for acquiring core knowledge in general. To insist otherwise would be to indicate an exception to the multiple realizability of cognitive functions (Bechtel & Mundale, 1999; Cao, 2022) and to claim that specific cognitive capacities cannot in principle emerge from domain-general learning—even when scaled up with sufficient data and computation (Long, 2024).
Another parallel concern along this line has been that pre-linguistic infants appear to lack explicit symbolic systems, and no widely accepted mechanisms explain how domain-specific conceptual representations could form solely through perceptual input (Carey, 2011). At the same time, however, reasoning abilities in LLMs appear to emerge entirely from linguistic generation mechanisms, even for capacities that do not depend on language in humans (Han et al., 2025). Thus, ruling out the possibility that LLM could have capacity to manipulate and abstract over linguistic symbols that enables the emergence of core knowledge representations through pure language learning would require strong theoretical justification (Long, 2024).
4.2 Groundless Learning Trajectory
Secondly, it could be noted that a central challenge in aligning LLM learning with human cognition lies in the temporal dynamics of data exposure. In human development, cognitive growth unfolds through a structured trajectory: infants begin with limited processing capacities and gradually acquire increasingly abstract representations, scaffolded by embodied interaction and core knowledge primitives (Pezzulo et al., 2013). This progression allows higher-order reasoning to emerge as a natural extension of early sensorimotor experiences. In contrast, LLMs are exposed to vast amounts of heterogeneous textual data all at once—without any temporal ordering, developmental constraints, or grounded scaffolding. As a result, they process low-level and high-level concepts simultaneously, often without forming the kind of layered, hierarchical conceptual structure that supports flexible and context-sensitive reasoning in humans (Mitchell & Krakauer, 2023). Recognizing this gap has motivated growing interest in curriculum learning, training strategies that introduce concepts in an incremental order from simple to complex, as a way to induce more structured, human-like learning trajectories in artificial systems (Liu et al., 2025).
However, the developmental mismatch in temporal sequencing should not be regarded as an inherent limitation of LLMs. The framework of grounded cognition does not prescribe a fixed developmental trajectory, rather, its central claim is that meaningful connections must be established between perceptual modalities and abstract reasoning. In this light, MLLMs are well-positioned to meet this criterion, by integrating visual, embodied, and linguistic inputs, they can construct grounded symbolic representations that exceed what is achievable through text-only training alone (Ma et al., 2024; Shi et al., 2025; Sheta et al., 2025). Developmentally structured data, organized to progress from simple to complex, may serve as a valuable complementary strategy in supporting this process. Curriculum learning approaches of this kind have been shown to enhance multimodal learning, especially when combined with large-scale text-based pretraining. This is particularly relevant for training core knowledge representations, which require the alignment of low-level sensorimotor regularities with higher-order symbolic abstraction. Nevertheless, the analogy to human development should not be overstretched, for instance, by assuming that models must be trained exclusively on data comparable to that available to children. LLMs and MLLMs rely on fundamentally different learning faculties, and the goal is not to imitate human developmental biology, but to implement functionally equivalent scaffolding that enables the emergence of structured, generalizable reasoning. From this perspective, the central priority is to foster robust links between multimodal representations and inferential capabilities, not to rigidly follow a human-like learning timeline.
@@ 4.3 Buried Too Deeply
The third, and perhaps most persuasive, interpretation of core knowledge deficits in MLLMs is that LLMs may, in fact, possess core knowledge but are unable to effectively extract and apply it during reasoning. Since the early days of connectionist research, it has been suggested that artificial neural networks, due to their reliance on distributed representations, inevitably face a challenge: as network size increases through scaling, retrieving knowledge representations for reasoning becomes increasingly difficult and computationally costly (Hinton et al., 1986; Clark, 1992). As Chalmers (1990) remarked:
@@
This concern is particularly salient in the context of core knowledge representations. High-level concepts required for complex tasks—such as recalling a specific historical event or recipe—are often encoded in dense, localized regions of the model’s representational space. Their distinctiveness and clustering make them relatively easy to retrieve. In contrast, low-dimensional concepts like perceptual constancy (i.e., the notion that an object’s identity persists despite changes in sensory input) are diffusely represented across many training examples with similar salience. Because foundational concepts like these emerge from broad contextual regularities rather than being tied to distinct lexical or visual anchors, language models often struggle to isolate and systematically apply them. Although such concepts are implicitly embedded throughout the training distribution, they are rarely reinforced in ways that facilitate robust retrieval or compositional reuse (Garrigan & Kellman, 2008; Green, 2024). This leads to brittle or shallow reasoning on tasks that require generalization from such principles. As a result, although increased scale may allow core knowledge to be encoded within the distributed weights of the model, this does not ensure that the knowledge is organized in a way that supports reliable inference, as the relevant representations often remain highly dispersed and entangled across the parameter space, hindering the model’s ability to access and deploy them for structured reasoning (Shani et al., 2023).
However, this interpretation also points to a potential opportunity for building core knowledge in LLMs. If such representations are already implicitly encoded within the model’s parameters, they may not be fundamentally absent but rather latent and insufficiently accessible. The key challenge, then, is not whether core knowledge can be represented in foundation models, but whether it can be activated employed effectively for reasoning. It may be thereby possible to exploit these distributed patterns through targeted training strategies, suggesting that the emergence of core knowledge in LLMs is not inherently unattainable.
@@ 5 Moving Forward: Growing AI Like A Child
Taken together, these interpretations suggest that while core knowledge may be innate in humans, it may not be necessarily unlearnable in artificial systems (Long, 2024). Innateness reflects evolved inductive biases, not architectural inevitabilities. Thus, core knowledge might be approximated in LLMs through targeted training that distills structured priors from curated, multi-modal data, making relevant principles salient in their knowledge representations. Rather than mimicking human biology, the goal is to functionally reconstruct core knowledge as a scaffold for structured, commonsense reasoning. For this proposal to be viable, we highlight several key conditions:
@@
    Sensory Grounding: models must be trained using multimodal inputs, incorporating both visual and linguistic data to approximate the rich perceptual experience of early human learning.
@@
    Large-scale Exposure: a sufficiently vast dataset is needed to ensure that core knowledge representations are robust and not lost due to extraction difficulties inherent in distributed neural representations.
@@
    Confound Minimization: unlike a child, LLMs are capable of leveraging advanced linguistic processing to access abstract information in their training data, making them prone to shortcut-taking by exploiting spurious correlations (Li et al., 2025b). Ensuring that training data is curated to prevent such biases is crucial.
To this end, this paper proposes a practical implementation strategy that satisfies the above criteria: large-scale pretraining on synthetic datasets generated via physics engines, curated through a novel cognitive prototyping strategy. The following section provides a detailed description of the approach and addresses potential concerns.
@@ 5.1 Cognitive Prototype
To generate training datasets that enhance the saliency of low-level cognitive representations while minimizing environmental confounds, one promising strategy is to directly train MLLMs on machine-adapted cognitive experiments widely used in developmental psychology. The recent translation of these paradigms into machine-readable formats—as exemplified by cognitively inspired benchmarks (Binz & Schulz, 2023; Li et al., 2025a)—provides a robust foundation for constructing such datasets.
@@
Many core cognitive concepts manifest in multiple forms and therefore require diverse experimental paradigms for robust assessment. To systematically structure the training process, we propose a standardized framework called cognitive prototyping—a methodology for curating data centered around individual cognitive abilities. A cognitive prototype includes a precise operationalization of a target concept, based on established cognitive tasks, along with schematic definitions of task-relevant conditions that support controlled variation and generalization. This approach enables large-scale data generation while preserving the conceptual rigor of developmental psychology experiments.
For example, the Three Mountain Task can serve as a prototype for training level-2 visual perspective-taking. In its most classic format, a child observes a three-dimensional model featuring three distinct mountains—one snow-covered, one marked with a red cross, and one topped with a hut—from multiple vantage points. The child is then asked to infer what a doll sees from a different viewpoint by selecting the correct image from a set of photographs (Piaget & Inhelder, 1969). This paradigm can be schematized into a range of task-relevant and task-irrelevant variables, such as the spatial positioning and visual appearance of the doll, the observer’s viewpoint offset, the model’s geometry, and the labeling configuration (Gao et al., 2024; Li et al., 2025a)  [^4] Li et al. (2025a) provides additional details on the concept prototyping procedure (see Fig.3). While said discussion focuses on curating evaluation sets, the same abstraction and control framework is equally applicable to constructing training datasets.. By systematically varying and recombining these parameters, large-scale and diverse training data can be generated while maintaining fidelity to the core conceptual structure of the task. This helps prevent overfitting to superficial cues and encourages models to abstract the underlying principle.
Importantly, cognitive prototypes also support cross-concept generalization within shared environments. For instance, the same Three Mountain setting, when paired with alternative labeling schemes or question types, can be adapted to train level-1 visual perspective-taking or mental rotation (Wang et al., 2025). This compositional structure allows hierarchically related cognitive abilities to be trained within a unified synthetic environment, enabling models to construct knowledge representations that reflect the interrelatedness of these concepts in the real world (Sun et al., 2024; Luo et al., 2024; Sun et al., 2025). By leveraging cognitive prototypes for structured data generation, this strategy offers a scalable and theoretically grounded approach to training MLLMs on core cognitive capacities. It addresses key limitations in current model development by combining large-scale learning with conceptually meaningful scaffolding.
@@ 5.2 Synthetic Data Via Physical Simulations
Building on the cognitive prototype framework, the next step is to generate large-scale datasets by systematically varying task conditions. Each generated instance represents a distinct combination of task parameters while preserving the underlying cognitive construct specified by the prototype. However, manually labeling such extensive datasets would be infeasible and would severely constrain both scale and diversity (Zhang et al., 2025). To overcome this limitation, we propose leveraging synthetic data generation—a computational strategy for producing richly annotated, scalable, and precisely controlled datasets aligned with cognitive paradigms.
Instead of relying on real-world data, task instances can be rendered using physics engines such as MuJoCo (Todorov et al., 2012) or Genesis (Zhou et al., ), which enable fine-grained manipulation of environmental and agent parameters. These simulation platforms offer the ability to instantiate cognitive experiments with high fidelity and reproducibility while systematically varying key conditions. By embedding task-relevant variables directly into the generation pipeline, such as object positions, agent viewpoints, labeling schemes, and visual occlusions, these engines can produce large volumes of synthetic data that maintain conceptual rigor while avoiding common artifacts and biases found in naturalistic datasets. In combination with cognitive prototypes, this simulation-based approach enables the scalable creation of high-quality training data for MLLMs. It promotes the development of core cognitive abilities by reinforcing structurally meaningful variations, discouraging shortcut learning, and scaffolding representations that mirror grounded knowledge representations in humans.
5.3 Addressing Concerns: can hollow representations be avoided?
A salient critique of synthetic training data is that it risks producing ”hollow representations”: knowledge representation that tailors only to highly rigid scenarios of core cognitive tasks, as opposed to grounded conceptual structures akin to human core knowledge that can be reliably leveraged in high-level reasoning. According to this critique, synthetic training environments, despite their controllability, lack the rich, embodied correspondence between perception and action that grounds human cognition. Consequently, the representations they produce might capture only surface-level statistical features, failing to generalize beyond their designed scope.
However, this concern rests on the assumption that grounding must emerge through human-like experiential data—an assumption we challenge under the “Buried Too Deeply” hypothesis introduced above. Core knowledge is not absent from large-scale pretraining corpora; rather, it is diffusely distributed across the representational manifold, making it difficult for models to access and reuse during reasoning. The limitation, therefore, lies not in the content of representations but in their retrievability and saliency. Knowledge distillation through synthetic data can address this challenge by providing cues for triggering associations between distributed representations through the scaffolding across simple and complex abilities along a given core knowledge domain, increasing their functional accessibility for reasoning.
This is supported under the standpoint of grounded cognition, robustness arises from grounded representational dependencies—where low-level sensory invariances serve as the substrate for higher-order abstraction. Crucially, this framework does not prescribe a fixed developmental trajectory. Temporal ordering in training—whether mimicking childhood or not—is secondary to associative accessibility: what matters is whether high-level reasoning reliably triggers the relevant low-level conceptual primitives when required. In this light, cognitively structured synthetic data do not aim to replicate human development but to restore associative coupling between representational layers that current MLLMs fail to exploit. Prototype-based synthetic data generation, following a structured hierarchical progression from simple to complex abilities, provides precisely this form of representational scaffolding. By encoding the latent structure of core cognitive concepts within parameterized and compositional synthetic environments, this approach enables the model to align distributed low-level representations with the higher-order inferential routines that depend on them. Such relational strengthening mitigates core knowledge dissociation—the tendency of MLLMs to exhibit symbolic reasoning that is divorced from grounded perception.
That said, our proposal addresses only one viable solution within the current MLLM paradigm. It remains possible, in principle, to construct models that acquire core knowledge through embodied exploration rather than data-driven reconstruction—effectively “raising” a model through interactive experience (Ahn et al., 2024a). Yet such an approach is unlikely to be scalable given the highly distributed, non-interactive nature of present foundation model infrastructure. Under existing architectures, large-scale synthetic data generation remains the most tractable means of making implicit world structure explicit in parameter space, enabling the emergence of robust, grounded representations without requiring human-like developmental trajectories.